# LW-FedSSL: Resource-Efficient Layer-wise Federated Self-Supervised Learning

### v2

@@ IV-A LW-FedMoCo
IV-B LW-FedSSL
To address the performance gap between layer-wise and end-to-end training for FedSSL, we propose LW-FedSSL by introducing two crucial mechanisms: server-side calibration, integrated at the server, and representation alignment, integrated into the local training process. Fig. 4 illustrates these two mechanisms of LW-FedSSL at stage $s$.
Server-side calibration: In layer-wise training, only the most recently added layer, $L_{s}$, is actively trained by clients at any given stage $s$. This may limit prior layers’ ability to adapt to the newly added layer, potentially degrading the model’s performance. Given a limited number of rounds $R_{s}$ available for each stage, solely training $L_{s}$ may prove insufficient to match the training received by the end-to-end strategy. On the other hand, increasing the number of rounds could introduce additional resource demands on clients. To tackle this, we can take advantage of resources available at the FL server. The server can contribute by training the current sub-model on a small auxiliary dataset, $D^{g}$, at the end of every communication round. Specifically, following the aggregation process, the server can train the aggregated encoder $F$ across all existing layers (i.e., $L_{1}$ to $L_{s}$) before broadcasting $F$ back to the clients. The auxiliary dataset $D^{g}$ may originate from a public dataset with a different data distribution than that of the clients’ data, as the server cannot directly access data within clients. Furthermore, $D^{g}$ may typically be unlabeled, requiring server-side training to also utilize an SSL approach. Training across all existing layers in the current sub-model plays a crucial role in calibrating the layers to function harmoniously. Also, our assumption regarding the auxiliary data at the server can be reasonably met, since in practice, the server may require a small dataset for various purposes, such as performance monitoring, pre-training, distillation, and other tasks in the FL context [23, 7, 24, 25, 26]. Particularly for FedSSL tasks, probing datasets may be necessary at the server to monitor and ensure smooth training progress.
Representation alignment: FL clients often exhibit distinct local data distributions resulting from diverse device usage patterns and application settings. Such data heterogeneity can introduce bias to the local training process, negatively affecting the performance of the global model. To mitigate this, some studies [27, 28] encourage local training processes across different clients to converge toward parameters that are closely aligned with the global parameters. Often, this is achieved by aligning the representations of local models with those of the global model while simultaneously pushing them away from old representations (from old local models of previous FL rounds) [27]. In contrast, our representation alignment mechanism does not require old local parameters. As a result, clients no longer need to store and perform inference on old parameters, reducing the overhead. Given a positive pair of views, $x_{q}$ and $x_{k^{+}}$, our approach uses local encoder $F^{i}$ and the global encoder $F$ to extract local representation $z^{i}=F^{i}(x_{q})$ and positive global representation $z^{+}=F(x_{k^{+}})$. Negative global representations $z^{-}$ are extracted from other samples. Similar to Equation 2, we can calculate the alignment loss as
The server-side calibration mechanism trains all existing layers in the global model to facilitate smooth collaboration between them. However, during local training, updates only to the active layer may disrupt this cohesion. To mitigate this, representation alignment encourages the updates of the local model to remain closely aligned with the parameters of the global model. This prevents complete overwriting and helps preserve the layer cohesion established by server-side calibration, while also allowing local models to incorporate local information from their respective local datasets. Additionally, representation alignment prevents extensive divergence between different local models, enabling a more effective FL aggregation process. In Section V-C, we empirically show that server-side calibration and representation alignment mechanisms are complementary for improving the effectiveness of layer-wise training. Integrating these mechanisms enables LW-FedSSL to compete with end-to-end training, all the while substantially reducing resource demands on FL clients. The details of LW-FedSSL are presented in Algorithms 1 and 2. Algorithm 1 outlines the server-side execution, which handles gradual increase in the model depth based on stage, server-side calibration mechanism, and aggregation. Each client executes Algorithm 2, performing local SSL training with representation alignment for regularization.
@@ IV-B LW-FedSSL
TABLE III: Linear evaluation and fine-tuning performance of different approaches. We conduct the experiments three times with different seeds. We compare the memory, FLOPs, and communication costs (download + upload) for a local training process. FedMoCo, FedBYOL, and FedSimCLR correspond to MoCo v3 [11], BYOL [29], and SimCLR [1] integrated with FL. For FedSimCLR, we use a batch size of 512. Three best values, except for the centralized MoCo v3, are marked in bold.
@@
IV-C Prog-FedSSL
Inspired by ProgFed in [18], we also explore progressive training for FedSSL, denoted as Prog-FedSSL. Similar to layer-wise training, progressive training also starts with an empty encoder $F$, appending a new layer $L_{s}$ to $F$ at each stage $s$. However, in contrast to updating only a single layer $L_{s}$, progressive training updates all existing layers (i.e., $L_{1}$ to $L_{s}$) during each stage $s$. This distinction results in higher computation and communication costs for Prog-FedSSL than layer-wise training approaches. Nonetheless, when compared with end-to-end training, which requires clients to train a full model with $S$ layers (i.e., $L_{1}$ to $L_{S}$) at each round, Prog-FedSSL enables clients to train only a sub-model with $s$ layers $(s\leq S)$. Moreover, clients only need to exchange the sub-model with the server. In essence, Prog-FedSSL saves both computation and communication costs for $(S-s)$ layers at each stage $s$ compared to end-to-end training. As we will demonstrate in Section V-B, progressive training outperforms both layer-wise and end-to-end training approaches. Nevertheless, it is important to note that the memory requirement of progressive training gradually increases as the stages progress, becoming equivalent to that of end-to-end training at the final stage (i.e., when $s=S$). This may limit low-memory devices from participating in the FL process, whereas the memory requirement of layer-wise training is significantly lower due to training a single layer $L_{s}$ at a time. We compare the different characteristics of FedMoCo (end-to-end), LW-FedMoCo, LW-FedSSL, and Prog-FedSSL in Table II. Fig. 15 in Appendix A illustrates FedMoCo, LW-FedMoCo, and Prog-FedSSL at stage $s$.
@@ IV-C Prog-FedSSL
Figure 6: Impact of batch size on different approaches. (a) Linear evaluation results on the STL-10 dataset and (b) peak memory requirements for different batch sizes.
V Experiment
V-A Experimental setup
[…]
Following common practices, we primarily use linear evaluation for performance comparison [36, 37, 1]. The projection head $H$ and prediction head $P$ are discarded, and a linear classifier is trained on top of the frozen encoder $F$ for 20 epochs using a batch size of 256. We use the AdamW optimizer with a base learning rate of 3e-2 and a weight decay of 1e-5. The learning rate decays with a cosine schedule. We use the RandAugment function available in the PyTorch [38] framework for augmentation. For fine-tuning, we set the base learning rate to 1e-3 and the number of epochs to 40 with a warmup period of 10 epochs. Other settings are similar to those of linear evaluation.
Figure 7: Linear evaluation performance of LW-FedSSL, the baseline LW-FedMoCo, and its two variants incorporating either server-side calibration or representation alignment. Weight transfer between blocks is not performed for these two variants, as they demonstrate better performance without it.
V-B Performance and Resource Comparison
@@
Table III compares the performance of different approaches. In terms of linear evaluation, LW-FedSSL demonstrates comparable performance to the end-to-end training of FedMoCo and significantly outperforms the purely layer-wise training of LW-FedMoCo. LW-FedSSL also outperforms or holds a competitive position against FedMoCo in fine-tuning. The progressive training approach, Prog-FedSSL, outperforms all other approaches in both linear evaluation and fine-tuning performance. Interestingly, the fine-tuning performance of FedMoCo is even lower than that of LW-FedMoCo in some cases. Please note that we apply the same fine-tuning settings, which may not be optimal for all approaches.
@@
In Fig. 5, we present resource requirements for a single client. Fig. 5a measures the memory usage at each local epoch, showing that the memory requirement of LW-FedSSL is $\times3$ lower than that of FedMoCo at peak usage. The memory usage of Prog-FedSSL gradually increases and eventually matches that of FedMoCo in the final stage. Fig. 5b shows that LW-FedSSL also maintains a lower FLOPs requirement than FedMoCo, with 2.1×2.1$\times$2.1 × lower total FLOPs consumption. Additional memory and FLOPs requirements of LW-FedSSL over LW-FedMoCo arise from the representation alignment mechanism, which requires inference on the global model to obtain global representations.
@@
In Figures 5c and 5d, we investigate the communication cost required for model exchange (download and upload) between the server and a client. Note that we only consider the encoder part for comparison, excluding the projection and prediction heads, which are common to all approaches with constant communication costs. In FedMoCo, a client needs to exchange the full encoder with the server at each communication round. Meanwhile, in LW-FedMoCo, a client only needs to exchange the active layer (i.e., $L_{s}$). Thus, both FedMoCo and LW-FedMoCo result in constant download and upload costs across different communication rounds. For Prog-FedSSL, both download and upload costs gradually increase as the number of active layers (i.e., $L_{1}$ to $L_{s}$) grows with the stage $s$. As for LW-FedSSL, a client needs to download layers $L_{1}$ to $L_{s}$ from the server, as these are updated through the server-side calibration mechanism. However, the client uploads a single layer $L_{s}$, as only it is updated by the local training process. Therefore, the download cost of LW-FedSSL increases with each step $s$, while the upload cost remains constant. Upon comparing the total communication cost over 180 rounds, LW-FedSSL is 1.8×1.8$\times$1.8 × cheaper in terms of download cost and 12×12$\times$12 × cheaper in terms of upload cost compared to FedMoCo.
@@
TABLE IV: Linear evaluation results of LW-FedSSL with different amounts of auxiliary data.
@@
V-C Ablation Study
In Fig. 7, we compare LW-FedSSL with the baseline LW-FedMoCo and its two variants, which solely incorporate either server-side calibration or representation alignment. Incorporating server-side calibration improves performance on the CIFAR-10 dataset (from which the auxiliary data is sampled), but leads to a decline on other datasets. This decline may result from conflicting update directions between training on auxiliary data and the local training. Compared to the baseline, incorporating representation alignment obtains better performance on the STL-10 dataset, with comparable performance on the others. Meanwhile, LW-FedSSL, which incorporates both mechanisms, consistently outperforms the baseline across all datasets. The server-side calibration mechanism facilitates smooth collaboration between different layers of the model. Meanwhile, representation alignment encourages the weights of the local model to remain closely aligned with those of the global model. This helps maintain layer cohesion established by the server-side calibration mechanism during the local training process. Fig. 7 suggests that instead of acting independently, server-side calibration and representation alignment are complementary for improving the performance of layer-wise training.
V-D Amount of Auxiliary Data
The server-side calibration mechanism of LW-FedSSL relies on the auxiliary dataset $D^{g}$. We vary the amount of auxiliary data at the server by sampling 1%, 10%, and 100% of the CIFAR-10 training data. We also conduct a centralized end-to-end SSL training using only the auxiliary data at the sever. The results in Table IV indicate that the performance of LW-FedSSL improves with an increase in the amount of data.
V-E Batch Size
SSL approaches tend to perform better with larger batch sizes due to a greater number of negative samples [1]. To investigate this, we set the local training batch size to 64, 128, 256, 512, and 1024. Fig. 6a indicates that layer-wise training approaches, LW-FedMoCo and LW-FedSSL, also tend to improve with increased batch size, similar to end-to-end SSL training. Fig. 6b reports peak memory requirements for each setting. Given that Prog-FedSSL does not freeze the layers, its peak memory requirement (at the last stage) is equivalent to that of FedMoCo. Memory requirements for FedMoCo and Prog-FedSSL sharply rise as the batch size grows, while those of LW-FedMoCo and LW-FedSSL remain relatively flat.
@@ V-E Batch Size
Figure 10: Number of participating clients in each FL round. The total number of clients is set to 45. Linear evaluation results are shown for the STL-10 dataset.
V-F Data Heterogeneity
@@
(a) Linear evaluation of LW-FedMoCo.
@@
Figure 13: Learning rate strategies.
V-G Total Number of Clients
To investigate the impact of the number of clients $N$, we conduct experiments with $N$ set to 5, 10, 30, and 45. For each setting, we uniformly partition the unlabeled data portion of the STL-10 dataset into $N$ local datasets. Therefore, the total number of data samples remains constant across all settings. Linear evaluation results with varying $N$ values are shown in Fig. 9. We can observe that all approaches, except LW-FedSSL, experience a drop in performance as the number of clients increases. Integration of representation alignment mechanism enables LW-FedSSL to maintain good performance even in scenarios with a large number of clients [27].
V-H Number of Participating Clients
@@
V-I Distribution of Number of Rounds Per Stage
@@
For LW-FedMoCo, the left-skewed setting results in the worst performance. In this setting, earlier layers receive updates for a small number of rounds, potentially lowering the quality of their output features. Even though later layers are trained for a large number of rounds, poor output features from the early layers may degrade the output of subsequent layers. The other two settings achieve comparable performance. As for LW-FedSSL, it allows early layers to continuously receive updates through the server-side calibration mechanism, which improves the performance of the left-skewed setting in most cases.
@@
In the case of Prog-FedSSL, none of the layers from prior stages are frozen, allowing a layer $L_{s}$ to receive training for all the rounds allocated through stages $s$ to $S$. This means that in a left-skewed allocation scenario, if we assign 10 rounds to stage $S-1$ and 50 rounds to stage $S$, then the layer $L_{S-1}$ undergoes training for a total of 60 rounds (since it is not frozen), while $L_{S}$ is trained for 50 rounds. However, if we reverse the scenario in the right-skewed allocation, i.e., assign 50 rounds to stage $S-1$ and 10 rounds to stage $S$, the layer $L_{S-1}$ still undergoes training for 60 rounds, while $L_{S}$ is trained only for 10 rounds. As a result, the left-skewed setting achieves better performance than the right-skewed setting; however, it would also consume more computation resources. The same case also applies to the server-side calibration mechanism of LW-FedSSL.
(a) Linear evaluation of LW-FedMoCo.
(b) Linear evaluation of LW-FedSSL.
(c) Linear evaluation of Prog-FedSSL.
Figure 14: Impact of different learning rate strategies.
V-J Learning Rate Strategy
By default, we use cosine decay scheduling for the learning rate. We explore different learning rate strategies for LW-FedMoCo, LW-FedSSL, and Prog-FedSSL, including a fixed learning rate and cosine decay scheduling for each training stage, which we refer to as “cyclic”. Fig. 13 illustrates how the learning rate changes at each local epoch for different strategies, and the results are presented in Fig. 14. For LW-FedMoCo, using a fixed learning rate achieves the best performance, and also for both LW-FedSSL and Prog-FedSSL, simply using a fixed learning rate can be a decent choice. A fixed learning rate may allow layers in later stages to receive effective updates, whereas in cosine decay scheduling, later layers undergo minimal updates due to a small learning rate. While the cyclic learning rate pattern proves to work decently for Prog-FedSSL, it may not be suitable for LW-FedMoCo and LW-FedSSL. As seen in Fig. 13, the learning rate for cyclic strategy drops quickly in each stage when only a limited number of rounds are available. This can impact layer-wise training, where a layer receives training only in its corresponding stage. (For LW-FedSSL with a fixed learning rate, we do not use weight transfer, as it demonstrates improved performance.)
VI Conclusion

### v3

@@ 4.2 LW-FedSSL
4.3 Prog-FedSSL
Inspired by ProgFed [24], we explore progressive training for FedSSL, denoted as Prog-FedSSL. Similar to layer-wise training, progressive training also starts with an empty encoder $F$, appending a new layer $L_{s}$ to $F$ at each stage $s$. However, in contrast to updating only a single layer $L_{s}$, progressive training updates all existing layers (i.e., $L_{1}$ to $L_{s}$) during each stage $s$. This distinction results in higher computation and communication costs for Prog-FedSSL than LW-FedSSL. Nonetheless, when compared with end-to-end training, which requires clients to train a full model with $S$ layers (i.e., $L_{1}$ to $L_{S}$) at each round, Prog-FedSSL enables clients to train only a sub-model with $s$ layers $(s\leq S)$. Moreover, clients only need to exchange the sub-model with the server. In essence, Prog-FedSSL saves both computation and communication costs for $(S-s)$ layers at each stage $s$ compared to end-to-end training. As we will demonstrate in Section 5.2, progressive training has the potential to outperform both layer-wise and end-to-end training approaches.
Nevertheless, it is important to recognize that the resource requirement of progressive training gradually increases as the stages progress, becoming equivalent to that of end-to-end training at the final stage (i.e., when $s=S$). This growing demand could prevent resource-constrained devices from participating in the training, raising similar limitations faced by end-to-end approach, leading to the loss of valuable data from these devices. Meanwhile, the resource requirement of layer-wise training can be significantly lower at any stage due to training a single layer $L_{s}$ at a time. We compare the different characteristics of FedMoCo (end-to-end), FedMoCo-LW (layer-wise), LW-FedSSL, and Prog-FedSSL in Table 2. In A, Fig. 15 illustrates FedMoCo, FedMoCo-LW, and Prog-FedSSL at stage $s$.
5 Experiment
5.1 Experimental setup
Unless otherwise stated, we use the following settings in our experiments.
Model. We use a ViT-Tiny [34] backbone with 12 transformer blocks (i.e., $S=12$) as encoder $F$. We use an input dimension of $32\times 32\times 3$ with a patch size of 4. The projection head $H$ and the prediction head $P$ are implemented as a 3-layer MLP and a 2-layer MLP, respectively. For both $H$ and $P$, the hidden layer dimensions are set to 4096 and the output dimension is set to 256.
[…]
Evaluation. Following common practices, we primarily use linear evaluation for performance comparison [37, 38, 1]. The projection head $H$ and prediction head $P$ are discarded, and a linear classifier is trained on top of the frozen encoder $F$ for 20 epochs using a batch size of 256. We use the AdamW optimizer with a base learning rate of 3e-2 and a weight decay of 1e-5. The learning rate decays with a cosine schedule. We use the RandAugment function available in the PyTorch [39] framework for augmentation. For fine-tuning, we set the base learning rate to 1e-3 and the number of epochs to 40 with a warmup period of 10 epochs. Other settings are similar to those of linear evaluation.
Datasets. We uniformly split the unlabeled data portion of the STL-10 dataset [40] to construct the local data of FL clients. To evaluate the performance, we use the CIFAR-10/100 [41], STL-10 (labeled portion), and Tiny ImageNet [42] datasets. For LW-FedSSL, we use only 10% of CIFAR-10 training data, excluding labels, as the auxiliary dataset $D^{g}$.
Baselines. In line with existing FedSSL studies [13, 20, 6], we establish end-to-end training baselines by integrating state-of-the-art SSL algorithms—MoCo v3 [11], BYOL [16], and SimCLR [1]—with FL [5] to create FedMoCo, FedBYOL, and FedSimCLR, respectively. We include FedMoCo-LW and FLL [25] with depth dropout (DD) as layer-wise training baselines. We set the depth dropout rate to 50%, as in [25]. For FedSimCLR and FLL, we use a batch size of 512. We also include MoCo v3 as the centralized upper bound. We mainly compare LW-FedSSL against FedMoCo and FedMoCo-LW, given that they are all based on the same SSL approach.
@@ 5.1 Experimental setup
Table 3: Linear evaluation and fine-tuning performance of different approaches. We conduct the experiments three times with different seeds. We compare max memory, total FLOPs, and communication costs (download + upload) for a client. Three best values, except for the centralized MoCo v3, are marked in bold.
@@
Figure 6: Impact of batch size on different approaches. (a) Linear evaluation results on the STL-10 dataset, and (b) peak memory requirements for different batch sizes.
5.2 Performance and Resource Comparison
@@
In Table 3, we compare the performance of different approaches. In terms of linear evaluation, LW-FedSSL demonstrates comparable performance to the end-to-end training of FedMoCo and significantly outperforms the layer-wise baseline FedMoCo-LW. LW-FedSSL also outperforms or holds a competitive position against FedMoCo in fine-tuning. The progressive training approach, Prog-FedSSL, outperforms all other approaches in both linear evaluation and fine-tuning performance. Interestingly, the fine-tuning performance of FedMoCo is even lower than that of FedMoCo-LW in some cases. Please note that we apply the same fine-tuning settings, which may not be optimal for all approaches.
@@
In Fig. 5, we present resource requirements for a single client. Fig. 5a measures the memory usage at each local epoch, showing that the memory requirement of LW-FedSSL is $\times3$ lower than that of FedMoCo at peak usage. The memory usage of Prog-FedSSL gradually increases and eventually matches that of FedMoCo in the final stage. Fig. 5b shows that LW-FedSSL also maintains a lower FLOPs requirement than FedMoCo, with 2.1×2.1$\times$2.1 × lower total FLOPs consumption. Additional memory and FLOPs requirements of LW-FedSSL over FedMoCo-LW arise from the representation alignment mechanism, which requires inference on the global model to obtain global representations.
@@
In Figures 5c and 5d, we investigate the communication cost required for model exchange (download and upload) between the server and a client. Note that we only consider the encoder part for comparison, excluding the projection and prediction heads, which are common to all approaches with constant communication costs. In FedMoCo, a client needs to exchange the full encoder with the server at each communication round. Meanwhile, in FedMoCo-LW, a client only needs to exchange the active layer (i.e., $L_{s}$). Thus, both FedMoCo and FedMoCo-LW result in constant download and upload costs across different communication rounds. For Prog-FedSSL, both download and upload costs gradually increase as the number of active layers (i.e., $L_{1}$ to $L_{s}$) grows with the stage $s$. As for LW-FedSSL, a client needs to download layers $L_{1}$ to $L_{s}$ from the server, as these are updated through the server-side calibration mechanism. However, the client uploads a single layer $L_{s}$, as only it is updated by the local training process. Therefore, the download cost of LW-FedSSL increases with each stage $s$, while the upload cost remains constant. Upon comparing the total communication cost over 180 rounds, LW-FedSSL is 1.8×1.8$\times$1.8 × cheaper in terms of download cost and 12×12$\times$12 × cheaper in terms of upload cost compared to FedMoCo.
It is important to recognize that we are considering the best-case scenario for the baseline FedMoCo-LW (and FLL+DD in Table 3), where each client participates in every training round. In this case, only the aggregated layer $L_{s}$ needs to be downloaded from the server, resulting in lower communication costs than LW-FedSSL. However, when new clients join or existing clients miss certain training stages, they need to download the current submodel ($L_{1}$ to $L_{s}$) from the server, similar to LW-FedSSL.
5.3 Ablation Study
Figure 7: Linear evaluation performance of FedMoCo-LW, LW-FedSSL, and its two variants, each incorporating either server-side calibration or representation alignment. Weight transfer between stages is not performed for these two variants, as they demonstrate better performance without it.
In Fig. 7, we compare the baseline FedMoCo-LW with LW-FedSSL and its two variants, which solely incorporate either server-side calibration or representation alignment. Incorporating server-side calibration improves performance on the CIFAR-10 dataset (from which the auxiliary data is sampled), but leads to a decline on other datasets. This decline may result from conflicting update directions between training on auxiliary data and the local datasets. Compared to the baseline, incorporating representation alignment obtains better performance on the STL-10 dataset, with comparable performance on the others. Meanwhile, LW-FedSSL, which incorporates both mechanisms, consistently outperforms the baseline across all datasets. The server-side calibration mechanism facilitates smooth collaboration between different layers of the model. Meanwhile, representation alignment helps maintain layer cohesion established by server-side calibration during the local training process. Fig. 7 suggests that instead of acting independently, server-side calibration and representation alignment are complementary for improving the performance of layer-wise training.
@@ 5.3 Ablation Study
Table 4: Linear evaluation results of LW-FedSSL with different amounts of auxiliary data.
@@
5.4 Amount of Auxiliary Data
LW-FedSSL relies on the auxiliary dataset $D^{g}$ for the server-side calibration mechanism. We vary the amount of auxiliary data at the server by sampling 1%, 10%, and 100% of the CIFAR-10 training data. We also conduct a centralized end-to-end SSL training using only $D^{g}$ from the sever. As expected, the results in Table 4 indicate that the performance of LW-FedSSL improves with an increase in the amount of data.
(a) Linear evaluation of FedMoCo-LW.
(b) Linear evaluation of LW-FedSSL.
(c) Linear evaluation of Prog-FedSSL.
Figure 8: Impact of different learning rate strategies.
5.5 Batch Size
SSL approaches tend to perform better with larger batch sizes due to a greater number of negative samples [1]. To investigate this, we set the local training batch size to 64, 128, 256, 512, and 1024. Fig. 6a indicates that layer-wise training approaches, FedMoCo-LW and LW-FedSSL, also tend to improve with increased batch size, similar to end-to-end SSL training. Fig. 6b reports peak memory requirements for each batch size setting. Given that Prog-FedSSL does not freeze the layers, its peak memory requirement (at the last stage $S$) is equivalent to that of FedMoCo. Memory requirements for FedMoCo and Prog-FedSSL sharply rise as the batch size grows, while those of FedMoCo-LW and LW-FedSSL remain relatively flat. When using comparable memory resources (as seen with the batch size of 64), LW-FedSSL outperforms FedMoCo while also retaining its layer-wise communication efficiency advantage.
@@ 5.5 Batch Size
5.6 Data Heterogeneity
@@
Figure 11: Number of participating clients in each FL round. Linear evaluation results are for STL-10.
Figure 12: Learning rate strategies.
5.7 Total Number of Clients
@@
(a) Linear evaluation of FedMoCo-LW.
@@
To investigate the impact of the number of clients $N$, we conduct experiments with $N$ set to 5, 10, 30, and 45. For each setting, we uniformly partition the unlabeled data portion of the STL-10 dataset into $N$ local datasets. Therefore, the total number of data samples remains constant across all settings. Linear evaluation results with varying $N$ values are shown in Fig. 10. We can observe that all approaches, except LW-FedSSL, experience a drop in performance as the number of clients increases. The representation alignment mechanism helps prevent significant divergence between local models, particularly in scenarios with a large number of clients [31, 33], enabling LW-FedSSL to maintain robust performance.
5.8 Number of Participating Clients
@@
5.9 Learning Rate Strategy
By default, we use cosine decay scheduling for the learning rate. We explore different learning rate strategies for FedMoCo-LW, LW-FedSSL, and Prog-FedSSL, including a fixed learning rate and cosine decay scheduling for each training stage, which we refer to as “cyclic”. Fig. 12 illustrates how the learning rate changes at each local epoch for different strategies, and the results are presented in Fig. 8. For FedMoCo-LW, using a fixed learning rate achieves the best performance, and also for both LW-FedSSL and Prog-FedSSL, simply using a fixed learning rate can be a decent choice. A fixed learning rate may allow layers in later stages to receive effective updates, whereas in cosine decay scheduling, later layers undergo minimal updates due to a small learning rate. While the cyclic learning rate pattern proves to work decently for Prog-FedSSL, it may not be suitable for FedMoCo-LW and LW-FedSSL. As seen in Fig. 12, the learning rate for cyclic strategy drops quickly in each stage when only a limited number of rounds are available. This can impact layer-wise training, where a layer receives training only in its corresponding stage.
5.10 Distribution of Number of Rounds Per Stage
@@
For FedMoCo-LW, the left-skewed setting results in the worst performance. In this setting, earlier layers receive updates for a small number of rounds, potentially lowering the quality of their output features. Even though later layers are trained for a large number of rounds, poor output features from the early layers may degrade the output of subsequent layers. The other two settings achieve comparable performance. As for LW-FedSSL, it allows early layers to continuously receive updates through the server-side calibration mechanism, which improves the performance of the left-skewed setting in most cases.
@@
In the case of Prog-FedSSL, none of the layers from prior stages are frozen, allowing layer $L_{s}$ to receive training for all the rounds allocated through stages $s$ to $S$. This means that in a left-skewed allocation scenario, if we assign 10 rounds to stage $(S-1)$ and 50 rounds to stage $S$, then the layer $L_{S-1}$ undergoes training for a total of 60 rounds (since it is not frozen), while $L_{S}$ is trained for 50 rounds. However, if we reverse the scenario in the right-skewed allocation, i.e., assign 50 rounds to stage $(S-1)$ and 10 rounds to stage $S$, the layer $L_{S-1}$ still undergoes training for 60 rounds, while $L_{S}$ is trained only for 10 rounds. As a result, the left-skewed setting achieves better performance than the right-skewed setting; however, it would also consume more resources. The same case also applies to the server-side calibration mechanism of LW-FedSSL.
6 Conclusion